# GraphTARIF: Linear Graph Transformer with Augmented Rank and Improved Focus

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph Neural Networks (GNNs) (Sperduti, 1993; Frasconi et al., 1998; Kipf and Welling, 2016; Hamilton et al., 2017; Velickovic et al., 2018; Luan et al., 2022) have emerged as a powerful paradigm for analyzing graph-structured data in a variety of Web-related domains, including social networks (Chen et al., 2017; Luan et al., 2019), recommender systems (Ying et al., 2018), communication networks (Lu et al., 2024a), and anomaly detection  (Dou et al., 2020). By leveraging message passing across graph neighborhoods, GNNs effectively capture local structural patterns and have achieved remarkable success in node classification, link prediction, and graph-level inference tasks. Nevertheless, their reliance on strictly local aggregation makes it inherently difficult to model long-range dependencies across distant nodes (Alon and Yahav, 2020), thereby restricting expressiveness and limiting their ability to approximate complex relational functions in large-scale Web graphs (Xu et al., 2018).
[…]
Inspired by recent advances in Transformer-based models in language and vision domains (Vaswani et al., 2017; Dosovitskiy et al., 2020), Graph Transformers (GTs)  (Ying et al., 2021; Ma et al., 2023) have gained increasing popularity in recent years. Unlike GNNs, GTs allow each node to directly attend to every other node in the graph, effectively creating a fully connected graph to mitigate locality bias and capture long-range interactions between distant nodes. The global self-attention (SA) mechanism can enhance the expressivity of the model and has demonstrated considerable success, especially in graph-level tasks such as molecular property prediction (Ma et al., 2023).
Nevertheless, the quadratic complexity of full self-attention ($O(N^{2})$ for $N$ nodes) poses a significant bottleneck for scaling Graph Transformers to node-level tasks on large-scale Web graphs, such as social networks or e-commerce interaction graphs, which often contain millions of nodes and edges. To address this challenge, recent works introduce linear self-attention (LSA) mechanisms (Wu et al., 2023, 2022; Deng et al., 2024), which approximate full softmax attention through kernel feature mappings or low-rank projections. This design reduces both memory and runtime costs from $O(N^{2})$ to $O(N+|E|)$, enabling GTs to process much larger graphs more efficiently. Despite these computational benefits, empirical studies consistently show that LSA-based Graph Transformers underperform compared to full-attention models or even classical GNNs across a variety of benchmarks (Luo et al., 2024; Zhou et al., 2024), raising concerns about their effectiveness in Web applications where both scalability and accuracy are critical.
@@ 2. Related Work
Graph Neural Networks. Graph Neural Networks (GNNs) have emerged as powerful methods for node classification tasks, primarily relying on local neighborhood aggregation through message passing. For example, Graph Convolutional Networks (GCNs) (Kipf and Welling, 2016) aggregate neighborhood information by leveraging a normalized adjacency matrix to propagate features. GraphSAGE (Hamilton et al., 2017) generalizes this approach by sampling and aggregating neighboring nodes, enabling inductive learning. Graph Attention Networks (GATs) (Velickovic et al., 2018) introduce attention mechanisms at the edge level, adaptively weighting neighbors based on their importance. Despite their success, these methods inherently suffer from local heterophily (Luan et al., 2024a, b; Lu et al., 2024b; Zheng et al., 2024, 2025), oversmoothing (Li et al., 2018; Luan et al., 2019), and oversquashing (Alon and Yahav, 2020; Topping et al., 2021), limiting their capability to represent complex node interactions across long-range distances.
@@ 3.1. Graph Neural Networks
where $h_{i}^{(l)}$ is the embedding of node $i$ at the $l$-th layer, $\mathcal{N}_{i}$ denotes the neighbors of node $i$, $f^{(l)}$ is the activation function, and $\text{AGG}^{(l)}(\cdot)$ is an aggregation function such as sum, mean, or max pooling (Xu et al., 2018; Hamilton, 2020).
@@ 4. Motivation
Furthermore, Zhang et al(Zhang et al., 2024) demonstrate that linear attention tends to produce smooth and high-entropy attention distributions. This issue becomes even more pronounced in graph-based tasks with a large number of nodes, where the attention weights are excessively dispersed. Such diffuse attention exacerbates the rank deficiency of the attention matrix, ultimately hindering the model’s expressiveness.
@@
Let ${\bm{X}}\in\mathbb{R}^{n\times d}$ be the node feature matrix and ${\bm{M}}\in\mathbb{R}^{n\times n}$ an attention matrix applied to transform the embeddings. Suppose that the rows of $\bm{X}$} are drawn from a Gaussian mixture model. Then, the expected inter-class variance after applying the attention transformation, measured by the trace of the between-class scatter matrix, is upper bounded by a term that scales linearly with the rank of $\bm{M}$}:
@@ 5.1. High-Rank Linear Attention
Overall, the proposed hybrid attention mechanism retains the computational efficiency of linear attention while incorporating an interpretable and controllable local inductive bias. This formulation effectively increases the rank of the attention score matrix and enhances the model’s representational capacity. Moreover, based on our experiments and prior studies, global graph attention often allocates excessive weights to distant nodes, which leads to overly smooth output representations and weakens the preservation of node-specific information.
@@
In summary, we propose a simple yet effective architectural that increases the rank of the attention scores matrix and reduces the entropy of the output representations, thereby enhancing the model’s capacity to produce discriminative node embeddings. Detailed visualizations can be found in Figs. 9, and  10 of Appendix C.
@@ 6. Experiments
6.1. Experimental Setup
6.1.1. Datasets
6.1.2. Baselines
We compare the proposed model against a broad set of baselines covering both classical GNNs and recent Graph Transformer architectures. Specifically, we include: (i) classical GNNs such as GCN, and GraphSAGE; (ii) heterophily-specialized GNNs, such as H2GCN, CPGNN, and TunedGNN, which is a recent framework that enhances traditional GNNs through skip-connections and hyperparameter optimization; and (iii) Graph Transformer models, such as GraphGPS, Exphormer, and NodeFormer. These baselines span a wide range of model paradigms and design philosophies, and collectively represent the current state-of-the-art under both homophilic and heterophilic settings. For baseline methods, we report the results as documented in prior work. For our method, we report the average performance over five independent runs.
6.2. Main Comparison
6.2.1. Results on Heterophilic Datasets
@@ 6.2.2. Results on Homophilic Datasets
As shown in Tab. 2, our model achieves the best or comparable performance across all five homophilic datasets. Our method achieves the highest average accuracy on the Computer (94.61%), Photo (97.03%), and CS (96.51%) datasets, and obtains the best average rank of 1.4 among all 19 baseline models. Compared to strong Graph Transformer baselines such as Polynormer, our model consistently improves performance by a clear margin, demonstrating its superior expressiveness on homophilic graphs. Considering that these baselines—such as the Tuned-GNN (Luo et al., 2024) models—have been carefully optimized on these mature benchmarks, the improvements achieved by our model are noteworthy and demonstrate its significant potential. These results validate that our model not only maintains high expressiveness on classical heterophilic graphs but also generalizes well to classical homophilic graphs.
@@ 6.2.3. Results on Large-scale Graphs
To further demonstrate the scalability of our approach on large-scale graphs, we evaluate GraphTARIF on three widely used benchmarks: ogbn-arxiv, pokec, and ogbn-products. Since the graphs in ogbn-products and pokec are too large to be loaded into GPU memory for full-batch training, we adopt the random partitioning strategy used in prior GT models (Wu et al., 2022, 2023) to perform mini-batch training. Among these benchmarks, ogbn-arxiv and ogbn-products are homophilic graphs, while pokec is heterophilic. As shown in Tab. 3, GraphTARIF achieves the best performance on ogbn-arxiv and on pokec. On ogbn-products, GraphTARIF also delivers highly competitive results. These results demonstrate that GraphTARIF can effectively scale to large-scale graphs and consistently achieves superior or competitive performance compared to strong baselines.
@@ 6.4.3. Efficiency Comparison
(a) Linear Attention
Rank=3, Entropy=9.37
(b) w/ $f(\cdot)$
Rank=190, Entropy=9.15
(c) w/ $f(\cdot)$ and Rank-Aug.
Rank=1000, Entropy=9.01
Figure 7. Visualization of attention matrices under three configurations on the Minesweeper dataset. Our method significantly increases the matrix rank while reducing the entropy.
6.4.4. Visualization of Attention Matrices
We visualize the attention matrices of the last layer on the Minesweeper dataset, as shown in Fig. 7. The vanilla linear attention produces low-rank, high-entropy matrices that limit expressiveness, while our method significantly increases the rank and reduces entropy.
@@ Appendix A Datasets and Experimental Details
Table 5 summarizes the statistics of all thirteen datasets used in our experiments, along with the corresponding major hyperparameter settings. We adopt GATConv as the base GNN operator preceding the attention layer, combined with residual connections, layer normalization, and ReLU activation. The attention module employs a linear attention mechanism with a sigmoid kernel feature map, enabling efficient global information aggregation. After the attention layer, one or two additional GNN layers are stacked to refine node representations and perform classification. All models are trained in a semi-supervised framework using the PyTorch Geometric (PyG) library. Experiments are conducted on a GPU cluster equipped with NVIDIA RTX 4090, RTX A6000, and L40S devices. Hyperparameters are optimized using Bayesian optimization via Optuna toolkit.
@@ B.1. Proof for Theorem 1
Although Theorem 1 provides a relatively loose upper bound on the inter-group variance, it still offers valuable insight in practice. Real-world data often exhibit complex, multi-modal, or manifold-structured distributions rather than simple class-wise Gaussian clusters. In such cases, a higher-rank attention matrix is still necessary to retain discriminative features across these intricate structures.
@@ B.2. Proof for Theorem 2
Here, we let ${\bm{M}}_{2}$ denote the conventional softmax-based attention matrix, and ${\bm{M}}_{1}$ denote the linear attention matrix derived from kernel-based or feature map approximations. We introduce a transformation operator $\bm{P}$} such that ${\bm{M}}_{1}={\bm{P}}{\bm{M}}_{2}$. Intuitively, $\bm{P}$} captures the structural bias introduced by linear attention, which tends to produce smoother, more homogeneous, and lower-rank attention distributions. To model this effect, it is natural to assume that $\bm{P}$} satisfies the energy contraction property, ${\bm{P}}^{\top}{\bm{P}}\preceq{\bm{I}}$, which reflects the smoothing and compression effects inherent to linear attention mechanisms.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/xxxxxxxxxxxxccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Neural networks
@@ 1. Introduction
Graph Neural Networks (GNNs) (Scarselli et al., 2008; Kipf, 2016; Hamilton et al., 2017; Veličković et al., 2017; Jian et al., 2025) have emerged as a powerful paradigm for analyzing graph-structured data in a variety of Web-related domains, including social networks (Chen et al., 2017), recommender systems (Ying et al., 2018), communication networks (Lu et al., 2024a), and anomaly detection  (Dou et al., 2020). By leveraging message passing across graph neighborhoods, GNNs effectively capture local structural patterns and have achieved remarkable success in node classification, link prediction, and graph-level inference tasks. Nevertheless, their reliance on strictly local aggregation makes it inherently difficult to model long-range dependencies across distant nodes (Alon and Yahav, 2020), thereby restricting expressiveness and limiting their ability to approximate complex relational functions in large-scale Web graphs (Xu et al., 2018).
Inspired by recent advances in Transformer-based models in language and vision domains (Vaswani et al., 2017; Dosovitskiy et al., 2021; Fan et al., 2022; Li et al., 2023; Chen et al., 2025), Graph Transformers (GTs)  (Ying et al., 2021; Ma et al., 2023) have gained increasing popularity in recent years. Unlike GNNs, GTs allow each node to directly attend to every other node in the graph, effectively creating a fully connected graph to mitigate locality bias and capture long-range interactions between distant nodes. The global self-attention (SA) mechanism can enhance the expressivity of the model and has demonstrated considerable success, especially in graph-level tasks such as molecular property prediction (Ma et al., 2023).
Nevertheless, the quadratic complexity of full self-attention ($O(N^{2})$ for $N$ nodes) poses a significant bottleneck for scaling Graph Transformers to node-level tasks on large-scale Web graphs, such as social networks or e-commerce interaction graphs, which often contain millions of nodes and edges. To address this challenge, recent works introduce linear self-attention (LSA) mechanisms (Wu et al., 2023, 2022; Deng et al., 2024), which approximate full softmax attention through kernel feature mappings or low-rank projections. This design reduces both memory and runtime costs from $O(N^{2})$ to $O(N+|E|)$, enabling GTs to process much larger graphs more efficiently. Despite these computational benefits, empirical studies show that LSA-based Graph Transformers underperform even classical GNNs across a variety of benchmarks (Luo et al., 2024; Zhou et al., 2024), raising concerns about their effectiveness in Web applications where both scalability and accuracy are critical.
@@ 2. Related Work
Graph Neural Networks. Graph Neural Networks (GNNs) have emerged as powerful methods for node classification tasks, primarily relying on local neighborhood aggregation through message passing. For example, Graph Convolutional Networks (GCNs) (Kipf, 2016) aggregate neighborhood information by leveraging a normalized adjacency matrix to propagate features. GraphSAGE (Hamilton et al., 2017) generalizes this approach by sampling and aggregating neighboring nodes, enabling inductive learning. Graph Attention Networks (GATs) (Veličković et al., 2017) introduce attention mechanisms at the edge level, adaptively weighting neighbors based on their importance. Despite their success, these methods inherently suffer from local heterophily (Luan et al., 2024a, b; Lu et al., 2024b; Zheng et al., 2024, 2025), oversmoothing (Li et al., 2018; Luan et al., 2019), and oversquashing (Alon and Yahav, 2020; Topping et al., 2021), limiting their capability to represent complex node interactions across long-range distances.
@@ 3.1. Graph Neural Networks
where $h_{i}^{(l)}$ is the embedding of node $i$ at the $l$-th layer, $\mathcal{N}_{i}$ denotes the neighbors of node $i$, $f^{(l)}$ is the activation function, and $\text{AGG}^{(l)}(\cdot)$ is an aggregation function such as sum, mean, or max pooling (Xu et al., 2018).
@@ 4. Motivation
Furthermore, Zhang et al. (Zhang et al., 2024) demonstrate that linear attention tends to produce smooth and high-entropy attention distributions. This issue becomes even more pronounced in graph-based tasks with a large number of nodes, where the attention weights are excessively dispersed. Such diffuse attention exacerbates the rank deficiency of the attention matrix, ultimately hindering the model’s expressiveness.
@@
Let ${\bm{X}}\in\mathbb{R}^{n\times d}$ be the node feature matrix and ${\bm{M}}\in\mathbb{R}^{n\times n}$ an attention matrix applied to transform the embeddings. Suppose that the rows of $\bm{X}$} are drawn from a Gaussian mixture model. The expected inter-class variance after applying the attention transformation, measured by the trace of the between-class scatter matrix, is upper bounded by a term that scales linearly with the rank of $\bm{M}$}:
@@ 5.1. High-Rank Linear Attention
Overall, the proposed hybrid attention mechanism retains the computational efficiency of linear attention while incorporating an interpretable and controllable local inductive bias. This formulation effectively increases the rank of the attention score matrix and enhances the model’s representational capacity. Moreover, based on our experiments and prior studies (Xing et al., 2024), global graph attention often allocates excessive weights to distant nodes, which leads to overly smooth output representations and weakens the preservation of node-specific information.
@@
In summary, we propose a simple yet effective architecture that increases the rank of the attention scores matrix and reduces the entropy of the output representations, thereby enhancing the model’s capacity to produce discriminative node embeddings. Detailed visualizations can be found in Figs. 8, and  9 of Appendix C.
@@ 6. Experiments
6.1. Experimental Setup
6.1.1. Datasets
6.1.2. Baselines
We compare the proposed model against a broad set of baselines covering both classical GNNs and recent Graph Transformer architectures. Specifically, we include: (i) classical GNNs such as GCN, and GraphSAGE. ; (ii) heterophily-specialized GNNs, such as H2GCN, CPGNN, and TunedGNN, which is a recent framework that enhances traditional GNNs through skip-connections and hyperparameter optimization. and (iii) Graph Transformer models, such as GraphGPS, Exphormer, and NodeFormer. These baselines span a wide range of model paradigms and design philosophies, and collectively represent the current state-of-the-art under both homophilic and heterophilic settings. For baseline methods, we report the results as documented in prior work. For our method, we report the average performance over five independent runs.
6.2. Main Comparison
6.2.1. Results on Heterophilic Datasets
@@ 6.2.2. Results on Homophilic Datasets
As shown in Tab. 2, our model achieves the best or comparable performance across all five homophilic datasets. Our method achieves the highest average accuracy on the Computer (94.61%), Photo (97.03%), and CS (96.51%) datasets, and obtains the best average rank of 1.4 among all 19 baseline models. Considering that these baselines—such as the Tuned-GNN (Luo et al., 2024) models—have been carefully optimized on these mature benchmarks, the improvements achieved by our model are noteworthy and demonstrate its significant potential. These results validate that our model not only maintains high expressiveness on classical heterophilic graphs but also generalizes well to classical homophilic graphs.
@@ 6.2.3. Results on Large-scale Graphs
To further demonstrate the scalability of our approach on large-scale graphs, we evaluate GraphTARIF on three widely used benchmarks: ogbn-arxiv, pokec, and ogbn-products. Among these benchmarks, ogbn-arxiv and ogbn-products are homophilic graphs, while pokec is heterophilic. As shown in Tab. 3, GraphTARIF achieves the best performance on ogbn-arxiv and on pokec. On ogbn-products, GraphTARIF also delivers highly competitive results. These results demonstrate that GraphTARIF can effectively scale to large-scale graphs and consistently achieves superior or competitive performance compared to strong baselines.
@@ 7. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (92570101) and the Fundamental Research Funds for the Central Universities (226-202500080).
@@ Appendix A Datasets and Experimental Details
Table  5 summarizes the statistics of all thirteen datasets used in our experiments, along with the corresponding major hyperparameter settings. We adopt GATConv as the base GNN operator preceding the attention layer, combined with residual connections, layer normalization, and ReLU activation. The attention module employs a linear attention mechanism with a sigmoid kernel feature map, enabling efficient global information aggregation. After the attention layer, one or two additional GNN layers are stacked to refine node representations and perform classification. All models are trained in a semi-supervised framework using the PyTorch Geometric (PyG) library. Experiments are conducted on a GPU cluster equipped with NVIDIA RTX 4090, RTX A6000, and L40S devices. Hyperparameters are optimized using Bayesian optimization via Optuna toolkit.
@@ B.1. Proof for Theorem 1
Although Theorem 1 provides a upper bound on the inter-group variance, it still offers valuable insight in practice. Real-world data often exhibit complex, multi-modal, or manifold-structured distributions rather than simple class-wise Gaussian clusters. In such cases, a higher-rank attention matrix is still necessary to retain discriminative features across these intricate structures.